Find the slope-intercept form for the line perpendicular to 2x - 3y = -6 and passing through (4, 9)


1
Expert's answer
2021-11-10T11:47:54-0500

The vector "(2,-3)" is a normal vector of the line "2x-3y=-6," and hence it is a collinear vector of the required line. Therefore, we get the equation of the line in the form "\frac{x-4}{2}=\frac{y-9}{-3}," which is equivalent to

"-3x+12=2y-18."

We conlcude that the slope-intercept form for the line perpendicular to 2x - 3y = -6 and passing through (4, 9) is

"y=-\frac{3}2x+15."